# CaRiNG: Learning Temporal Causal Representation under Non-Invertible Generation Process

### v1

@@ A.1 Proof for Theorem 1
Note that in the proof of Theorem 2, we require the transition lag $\tau$ to be larger than the mixing lag $r=1$. When a mixing lag exists, the guarantee of identifiability requires dynamic information from a further previous time step. As long as this inequality $\tau>r$ is satisfied, the parameters $\tau$ can be extended to arbitrary numbers following a similar modification in Appendix A.2.
A.2 Extension to Multiple Lags
@@
Multiple Transition Time Lag $\tau$. For the sake of simplicity, we consider only one special case with $\tau=r+1$ in Theorem 2. Our identifiability theorem can be actually extended to arbitrary lags directly. For any given $\tau$, according to modularity we have different conclusion at Eq 18 as $\textbf{LHS}=P(\mathbf{z}_{t}|\mathbf{x}_{t-1:t-\mu-r-\tau})=P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-\tau}).$ Similarity $\textbf{RHS}=P(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r-\tau})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-r-\tau})$ holds true as well. In addition, some modifications are needed in sufficiency assumption, i.e., re-define $\eta_{kt}\triangleq\log p(z_{kt}|\mathbf{z}_{t-1:t-r-\tau})$ and there should be at least $2n$ linear independent vectors for $\mathbf{v},\mathring{\mathbf{v}}$ with regard to $z_{l,\eta}$ where $l=1,2,\cdots,n$ and $t-\tau\leq\mu\leq t-1$. No extra changes are needed.
@@
Infinite Mixing Lag $r$. Theorem 1 can also be easily extended to infinite mixing lag since $\mathbf{\hat{z}}_{t}=\mathbf{h}(\mathbf{z}_{t};\mathbf{x}_{<t})$ still exists when $r\rightarrow\infty$, and the theorem still holds true.
A.3 Continuity for Permutation Invariance
@@
Let us first give an extreme example to illustrate the importance of extra constraints for identifiability when the lebesgue measure of $\mathbf{z}_{t}$ is not non-zero everywhere in $\mathbb{R}^{n}$. Consider 4 independent random variables $u,v,x,y$ subjects to standard normal distribution respectively. Suppose that there exist an invertible function $(x,y)=\mathbf{h}(u,v)$ satisfies
$\begin{cases}x=\mathbb{I}(x+y>0)\cdot u+\mathbb{I}(x+y\leq 0)\cdot v\\ y=\mathbb{I}(x+y>0)\cdot v+\mathbb{I}(x+y\leq 0)\cdot u.\end{cases}$ (24)
@@
In the literature of nonlinear ICA, the gap between $\mathbf{H}_{ij}\cdot\mathbf{H}_{ik}=0$ when $j\not=k$ and identifiability is ill-discussed. In linear ICA, since the Jacobian is a constant matrix, these two statements are equivalent. Nevertheless, in nonlinear ICA, $\mathbf{H}=\frac{\partial\mathbf{z}}{\partial\mathbf{\hat{z}}}$ is not a constant, but a function of $\mathbf{\hat{z}}$, which may leads to the failure of identifiability as shown in Eq 24.
@@
$\begin{cases}x=f(u,v)\cdot u+(1-f(u,v))\cdot v\\ y=f(u,v)\cdot v+(1-f(u,v))\cdot u.\end{cases}$ (25)
@@
Lemma 3 (Disentanglement with Continuity).
@@
${\left[\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\bigg{|}_{\hat{\mathbf{z}}=\mathbf{a}}\right]_{j}\not=0,\quad\left[\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\bigg{|}_{\hat{\mathbf{z}}=\mathbf{b}}\right]_{k}\not=0}$ (26)
where $[\cdot]_{j}$ refers to the $j$-th entry of vector. Their values are respectively within $\left\{(0,0,...,\frac{\partial{z}_{i}}{\partial{\hat{z}}_{j}},0,...,0)\in\mathbb{R}^{n}:\frac{\partial{z}_{i}}{\partial{\hat{z}}_{j}}\neq 0\right\}$ and $\left\{(0,0,...,\frac{\partial{z}_{i}}{\partial{\hat{z}}_{k}},0,...,0)\in\mathbb{R}^{n}:\frac{\partial{z}_{i}}{\partial{\hat{z}}_{k}}\neq 0\right\}$. Clearly, there is no path from $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{a}}$ to $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{b}}$. Since $\mathbf{h}$ is a second-order differentiable invertible function, we have its derivative $\mathbf{h}^{\prime}$ is also differentiable. Thus, $\mathcal{\hat{Z}}\subseteq\mathbb{R}^{n}$ is a path-connected domain which denotes that the image of $\frac{\partial{z}_{i}}{\partial\mathbf{\hat{z}}}$ is also path-connected. It will be violated that there is no path from $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{a}}$ to $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{b}}$ thus the proof is established.
@@
Let’s further review the example we provided earlier. Example in Eq 24 and Eq 25 respectively demonstrate the scenarios where the assumptions of differentiability and connectivity fail, leading to the breakdown of identifiability.
Lemma 4 (Disentanglement with Continuity under Side Information).
@@
${\left[\frac{\partial z_{i}}{\partial(\hat{\mathbf{z}},\mathbf{c})}\bigg{|}_{(\hat{\mathbf{z}},\mathbf{c})=\mathbf{a}}\right]_{j}\not=0,\quad\left[\frac{\partial z_{i}}{\partial(\hat{\mathbf{z}},\mathbf{c})}\bigg{|}_{(\hat{\mathbf{z}},\mathbf{c})=\mathbf{b}}\right]_{k}\not=0.}$ (27)
@@
Similar to Lemma 3, there exists no path between them because they are blocked in $\mathcal{\hat{Z}}$ alone. In the same way, since $\mathbf{h}$ is a second-order differentiable invertible function, and the non-degeneracy condition holds, the image of $\frac{\partial z_{i}}{\partial(\mathbf{\hat{z}},\mathbf{c})}$ is also path-connected. It will be violated and the proof is established.
A.4 Identifiability Benefits from Non-Stationarity
@@
We can further leverage the advantage of non-stationary data for identifiability. Let $\mathbf{v}_{kt}(u_{r})$ be $\mathbf{v}_{kt}$, which is defined in Eq 12, in the $u_{r}$ context. Similarly, Let $\mathring{\mathbf{v}}_{kt}(u_{r})$ be $\mathring{\mathbf{v}}_{kt}$ in the $u_{r}$ context. Let
$\mathbf{s}_{kt}\triangleq\Big{(}\mathbf{v}_{kt}(u_{1})^{\intercal},...,\mathbf{v}_{kt}(u_{m})^{\intercal},\frac{\partial^{2}\eta_{kt}({u}_{2})}{\partial z_{kt}^{2}}-\frac{\partial^{2}\eta_{kt}({u}_{1})}{\partial z_{kt}^{2}},...,\frac{\partial^{2}\eta_{kt}({u}_{m})}{\partial z_{kt}^{2}}-\frac{\partial^{2}\eta_{kt}({u}_{m-1})}{\partial z_{kt}^{2}}\Big{)}^{\intercal},$
$\mathring{\mathbf{s}}_{kt}\triangleq\Big{(}\mathring{\mathbf{v}}_{kt}(u_{1})^{\intercal},...,\mathring{\mathbf{v}}_{kt}(u_{m})^{\intercal},\frac{\partial\eta_{kt}({u}_{2})}{\partial z_{kt}}-\frac{\partial\eta_{kt}({u}_{1})}{\partial z_{kt}},...,\frac{\partial\eta_{kt}({u}_{m})}{\partial z_{kt}}-\frac{\partial\eta_{kt}({u}_{m-1})}{\partial z_{kt}}\Big{)}^{\intercal}.$
@@
As provided below, in our case, the identifiability of $\mathbf{z}_{t}$ is guaranteed by the linear independence of the whole function vectors $\mathbf{s}_{kt}$ and $\mathring{\mathbf{s}}_{kt}$, with $k=1,2,...,n$. This linear independence is generally a much stronger condition.
Corollary 1 (Identifiability under Non-Stationary Process).
@@
Suppose $\mathbf{x}_{t}=\mathbf{g}(\mathbf{z}_{t})$, $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$, and that the conditional distribution $p(z_{k,t}\,|\,\mathbf{z}_{t-1},\mathbf{u})$ may change across $m$ values of the context variable $\mathbf{u}$, denoted by $u_{1}$, $u_{2}$, …, $u_{m}$. Suppose the components of $\mathbf{z}_{t}$ are mutually independent conditional on $\mathbf{z}_{t-1}$ in each context. Assume that the components of $\hat{\mathbf{z}}_{t}$ are also mutually independent conditional on $\hat{\mathbf{z}}_{t-1}$. Suppose the domain is path-connected and $\mathbf{m},\mathbf{\hat{m}},\mathbf{g},\mathbf{\hat{g}}$ are second-order differentiable and their combination subjects to non-degenerate condition. If the $2n$ function vectors $\mathbf{s}_{k,t}$ and $\mathring{\mathbf{s}}_{k,t}$, with $k=1,2,...,n$, are linearly independent, then $\hat{\mathbf{z}}_{t}$ is a permuted invertible component-wise transformation of $\mathbf{z}_{t}$.

### v2

@@ Acknowledgments
Note that in the proof of Theorem A1, we require the transition lag $\tau$ to be larger than the mixing lag $r=1$. When a mixing lag exists, the guarantee of identifiability requires dynamic information from a more previous time step. As long as this inequality $\tau>r$ is satisfied, the parameters $\tau$ can be extended to arbitrary numbers following a similar modification in Appendix A1.4.
A1.2 Discussion for the sufficiency assumption.
This assumption describes the changability of latent variables. Taking the video understanding as an example, the latent variables may represent the concepts. The linear independence of the latent variables means that there exists a characteristic of the concept that cannot be linearly represented by others. To further illustrate the sufficiency assumption, we give 2 examples (Yao et al., 2022a) to show when and when not the sufficiency assumption holds.
One possible distribution that breaks this assumption is the additive Gaussian noise. Denote $\mathbf{z}_{h}$ as historical parents. Let $z_{kt}=q_{k}(\mathbf{z}_{h})+\epsilon_{kt}$ where $\epsilon_{kt}\sim N(0,1)$. In this case, we have $\eta_{kt}=\log P(z_{kt}|\mathbf{z}_{h})=-\log\sqrt{2\pi}-\frac{(z_{kt}-q_{k}(\mathbf{z}_{h}))^{2}}{2}$, and $\frac{\partial^{2}\log P(z_{kt}|\mathbf{z}_{h})}{\partial^{2}z_{kt}}=0$, which will violate the assumption.
On the opposite, if $\epsilon_{kt}$ subjects a zero mean generalized normal distribution: $P(\epsilon_{kt})\propto e^{-\lambda|\epsilon_{kt}|^{\beta}}$ with $\lambda>0$ and $\beta>2$ and $\beta\not=3$. Let $z_{kt}=q_{k}(\mathbf{z}_{h})+\epsilon_{kt}$ in which $q$ is a linear function. If for each $z_{kt}$ there exists at least one $k^{\prime}$ such that $c_{kk^{\prime}}=\frac{\partial z_{kt}}{\partial z_{k^{\prime},t-1}}\not=0$, the sufficiency assumption must hold.
In this case, we have
$\frac{\partial^{3}\eta_{kt}}{\partial^{2}z_{kt}\partial z_{k^{\prime},t-1}}=-\lambda\text{ sgn}(\epsilon_{kt})\beta(\beta-1)(\beta-2)|\epsilon_{kt}|^{\beta-3}c_{kk^{\prime}}$ (A13)
and
$\frac{\partial^{2}\eta_{kt}}{\partial z_{kt}\partial z_{k^{\prime},t-1}}=-\lambda\beta(\beta-1)|\epsilon_{kt}|^{\beta-2}c_{kk^{\prime}}.$ (A14)
We know that $|\epsilon_{lt}|^{\beta-2}$ and $|\epsilon_{lt}|^{\beta-2}$ are linearly independent since their ratio $|\epsilon_{lt}|$ is not constant. Besides, $|\epsilon_{lt}|^{\beta-2}$ and $|\epsilon_{lt}|^{\beta-2}$, with $l=1,2,\cdots,n$ are $2n$ linearly independent functions because of the different arguments involved. Suppose there exists $\alpha_{l1},\alpha_{l2}$ for $l=1,2,\cdots,n$, such that the weighted sum with regard to $\mathbf{v}_{l,t}$ is zero. Thus, for any $k$ we have
$\alpha_{k1}c_{kk^{\prime}}|\epsilon_{kt}|^{\beta-2}+\alpha_{k2}c_{kk^{\prime}}|\epsilon_{kt}|^{\beta-3}+\sum_{l\not=k}(\alpha_{l1}c_{lk^{\prime}}|\epsilon_{lt}|^{\beta-2}+\alpha_{l2}c_{lk^{\prime}}|\epsilon_{lt}|^{\beta-3})=0.$ (A15)
Since $|\epsilon_{kt}|^{\beta-2}$ and $|\epsilon_{kt}|^{\beta-3}$ with $l=1,2,\cdots,n$ are linearly independent and $c_{kk^{\prime}}\not=0$, the make the above equation holds, we have $\alpha_{k1}=\alpha_{k2}=0$. As this applies to any k, we know that $\alpha_{l1}$ and $\alpha_{l2}$ must be 0, for all $l=1,2,\cdots,n$. That is, $\{\mathbf{v}_{lt}\}$ is linearly independent. Thus, the sufficiency assumption holds.
Please note that the sufficiency assumption is crucial to the identifiability theory, yet not that restrictive. Even if it is not completely satisfied, we can still obtain some subspace identifiability (Kong et al., 2022).
A1.3 Discussion for the cross-time disentanglement
This section demonstrates how the entanglements between variables across time steps are prevented. Generally speaking, if the information is lost in the transition from $\mathbf{z}_{t}$ to $\mathbf{x}_{t}$, we have to borrow the information from context such as $\mathbf{z}_{t-1}$ to recover it. It is natural to receive information from $\mathbf{z}_{t-1}$ in order to find the best estimator.
Specifically, let us consider a generating process $x_{it}=g_{i}(\mathbf{z}_{t}),z_{it}=f_{i}(\mathbf{z}_{t-1},\epsilon_{it})$, where $\mathbf{x}_{t}\in R^{d},\mathbf{z}_{t}\in R^{n}$. Since $x_{it}$ can be fully charactized by $\mathbf{z}_{t}$, but not a function of $\mathbf{z}_{t-1}$, we have $\frac{\partial x_{i,t}}{\partial\mathbf{z}_{t-1}}=0$. For the estimation process $x_{i,t}=\hat{g}_{i}(\mathbf{\hat{z}}_{t})$, we have $\frac{\partial x_{i,t}}{\partial\mathbf{z}_{t-1}}=\frac{\partial x_{i,t}}{\partial\mathbf{\hat{z}}_{t}}\cdot\frac{\partial\mathbf{\hat{z}}_{t}}{\partial\mathbf{z}_{t-1}}=0$ for $i=1,2,\cdots,d$. Formally, we have an equation as $\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{\hat{z}}_{t}}\cdot\frac{\partial\mathbf{\hat{z}}_{t}}{\partial\mathbf{z}_{t-1}}=0$, i.e.,
$\begin{bmatrix}\frac{\partial x_{1,t}}{\partial\hat{z}_{1t}}&\cdots&\frac{\partial x_{1,t}}{\partial\hat{z}_{nt}}\\ \vdots&\ddots&\vdots\\ \frac{\partial x_{d,t}}{\partial\hat{z}_{1t}}&\cdots&\frac{\partial x_{d,t}}{\partial\hat{z}_{nt}}\\ \end{bmatrix}\cdot\begin{bmatrix}\frac{\partial\hat{z}_{1t}}{\partial z_{1,t-1}}&\cdots&\frac{\partial\hat{z}_{1t}}{\partial z_{n,t-1}}\\ \vdots&\ddots&\vdots\\ \frac{\partial\hat{z}_{nt}}{\partial z_{1,t-1}}&\cdots&\frac{\partial\hat{z}_{nt}}{\partial z_{n,t-1}}\\ \end{bmatrix}=\mathbf{0}.$ (A16)
If there exists at least $n$ different $i$ such that derivative of $x_{it}$ with respect to $\hat{\mathbf{z}}_{t}$ as $n$ vector functions $\begin{bmatrix}\frac{\partial x_{i,t}}{\partial\hat{z}_{1t}}&\cdots&\frac{\partial x_{i,t}}{\partial\hat{z}_{nt}}\end{bmatrix}$ are linearly independent, we have
$\begin{bmatrix}\frac{\partial\hat{z}_{1t}}{\partial z_{1,t-1}}&\cdots&\frac{\partial\hat{z}_{1t}}{\partial z_{n,t-1}}\\ \vdots&\ddots&\vdots\\ \frac{\partial\hat{z}_{nt}}{\partial z_{1,t-1}}&\cdots&\frac{\partial\hat{z}_{nt}}{\partial z_{n,t-1}}\\ \end{bmatrix}=\mathbf{0}$ (A17)
holds true. If the rank $R(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{\hat{z}}_{t}})$ of the matrix of $\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{\hat{z}}_{t}}$ is less than $n$, we have that the entanglement can only happened on the rest $n-R(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{\hat{z}}_{t}})$ dimensions.
Let us consider 2 extreme cases. First, if mixing function $\mathbf{g}$ is invertible, we have $R(\frac{\partial\mathbf{x}_{t}}{\partial\mathbf{\hat{z}}_{t}})=n$, and the entanglement between time steps are prevented. As another extreme case, in the NG setting as mentioned in Appendix A2.1, we set one dimension of the latent variable that is totally lost during the mixing process. In this case, we have to use $\mathbf{z}_{t-1}$ for its estimation. Even in this case, the unnecessary entanglements are prevented as well.
A1.4 Extension to Multiple Lags
@@
Multiple Transition Time Lag $\tau$. For the sake of simplicity, we consider only one special case with $\tau=r+1$ in Theorem A1. Our identifiability theorem can be actually extended to arbitrary lags directly. For any given $\tau$, according to modularity, we have different conclusion at Eq A7 as $\textbf{LHS}=P(\mathbf{z}_{t}|\mathbf{x}_{t-1:t-\mu-r-\tau})=P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-\tau}).$ Similarity $\textbf{RHS}=P(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r-\tau})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-r-\tau})$ holds true as well. In addition, some modifications are needed in sufficiency assumption, i.e., re-define $\eta_{kt}\triangleq\log p(z_{kt}|\mathbf{z}_{t-1:t-r-\tau})$ and there should be at least $2n$ linear independent vectors for $\mathbf{v}$ with regard to $z_{lt^{\prime}}$ where $l=1,2,\cdots,n$ and $t-\tau\leq t^{\prime}\leq t-r-1$. No extra changes are needed.
@@
Infinite Mixing Lag $r$. Theorem A1 can also be easily extended to infinite mixing lag since $\mathbf{\hat{z}}_{t}=\mathbf{h}(\mathbf{z}_{t};\mathbf{x}_{<t})$ still exists when $r\rightarrow\infty$, and the theorem still holds true.
A1.5 Continuity for Permutation Invariance
@@
Let us first give an extreme example to illustrate the importance of extra constraints for identifiability when the probability density of $\mathbf{z}_{t}$ is not non-zero everywhere in $\mathbb{R}^{n}$. Consider 4 independent random variables $u,v,x,y$ subject to standard normal distribution respectively. Suppose that there exists an invertible function $(x,y)=\mathbf{h}(u,v)$ satisfies
$\begin{cases}x=\mathbb{I}(x+y>0)\cdot u+\mathbb{I}(x+y\leq 0)\cdot v\\ y=\mathbb{I}(x+y>0)\cdot v+\mathbb{I}(x+y\leq 0)\cdot u.\end{cases}$ (A18)
@@
In the literature of nonlinear ICA, the gap between $\mathbf{H}_{ij}\cdot\mathbf{H}_{ik}=0$ when $j\not=k$ and identifiability is ill-discussed. In linear ICA, since the Jacobian is a constant matrix, these two statements are equivalent. Nevertheless, in nonlinear ICA, $\mathbf{H}=\frac{\partial\mathbf{z}}{\partial\mathbf{\hat{z}}}$ is not a constant, but a function of $\mathbf{\hat{z}}$, which may leads to the failure of identifiability as shown in Eq A18.
@@
$\begin{cases}x=f(u,v)\cdot u+(1-f(u,v))\cdot v\\ y=f(u,v)\cdot v+(1-f(u,v))\cdot u.\end{cases}$ (A19)
@@
Lemma A1 (Disentanglement with Continuity).
@@
${\left[\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\bigg{|}_{\hat{\mathbf{z}}=\mathbf{a}}\right]_{j}\not=0,\quad\left[\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\bigg{|}_{\hat{\mathbf{z}}=\mathbf{b}}\right]_{k}\not=0}$ (A20)
where $[\cdot]_{j}$ refers to the $j$-th entry of vector. Their values are respectively within $\left\{(0,0,...,\frac{\partial{z}_{i}}{\partial{\hat{z}}_{j}},0,...,0)\in\mathbb{R}^{n}:\frac{\partial{z}_{i}}{\partial{\hat{z}}_{j}}\neq 0\right\}$ and $\left\{(0,0,...,\frac{\partial{z}_{i}}{\partial{\hat{z}}_{k}},0,...,0)\in\mathbb{R}^{n}:\frac{\partial{z}_{i}}{\partial{\hat{z}}_{k}}\neq 0\right\}$. Clearly, there is no path from $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{a}}$ to $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{b}}$. Since $\mathbf{h}$ is a second-order differentiable invertible function, we have its derivative $\mathbf{h}^{\prime}$ is also differentiable. Thus, $\mathcal{\hat{Z}}\subseteq\mathbb{R}^{n}$ is a path-connected domain which denotes that the image of $\frac{\partial{z}_{i}}{\partial\mathbf{\hat{z}}}$ is also path-connected. It will be violated that there is no path from $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{a}}$ to $\frac{\partial z_{i}}{\partial\hat{\mathbf{z}}}\big{|}_{\hat{\mathbf{z}}=\mathbf{b}}$ thus the proof is established. ∎
@@
Let’s further review the example we provided earlier. Examples in Eq A18 and Eq A19 respectively demonstrate the scenarios where the assumptions of differentiability and connectivity fail, leading to the breakdown of identifiability.
Lemma A2 (Disentanglement with Continuity under Side Information).
@@
${\left[\frac{\partial z_{i}}{\partial(\hat{\mathbf{z}},\mathbf{c})}\bigg{|}_{(\hat{\mathbf{z}},\mathbf{c})=\mathbf{a}}\right]_{j}\not=0,\quad\left[\frac{\partial z_{i}}{\partial(\hat{\mathbf{z}},\mathbf{c})}\bigg{|}_{(\hat{\mathbf{z}},\mathbf{c})=\mathbf{b}}\right]_{k}\not=0.}$ (A21)
@@
Similar to Lemma A1, there exists no path between them because they are blocked in $\mathcal{\hat{Z}}$ alone. In the same way, since $\mathbf{h}$ is a second-order differentiable invertible function, and the non-degeneracy condition holds, the image of $\frac{\partial z_{i}}{\partial(\mathbf{\hat{z}},\mathbf{c})}$ is also path-connected. It will be violated and the proof is established.
A1.6 Identifiability Benefits from Non-Stationarity
@@
We can further leverage the advantage of non-stationary data for identifiability. We rewrite $\mathbf{v}_{lt}$, which is defined in Eq A3, as $\mathbf{s}_{lt}(u_{r})$ in the $u_{r}$ context as
$\mathbf{s}_{lt}(u_{r})\triangleq\Big{(}\frac{\partial^{2}\eta_{1t}(u_{r})}{\partial z_{1t}\partial z_{l,t-r-1}},...,\frac{\partial^{2}\eta_{nt}(u_{r})}{\partial z_{nt}\partial z_{l,t-r-1}},\frac{\partial^{3}\eta_{1t}(u_{r})}{\partial z_{1t}^{2}\partial z_{l,t-r-1}},...,\frac{\partial^{3}\eta_{nt}(u_{r})}{\partial z_{nt}^{2}\partial z_{l,t-r-1}}\Big{)}^{\intercal}.$ (A22)
We also consider the version of subtraction $\mathring{\mathbf{s}}_{t}(u_{r})$ from $u_{r}$ to $u_{0}$ without taking the derivative with respect to $z_{l,t-r-1}$ as
$\mathring{\mathbf{s}}_{t}(u_{r})\triangleq\Big{(}\frac{\partial^{2}\eta_{1t}(u_{r})}{\partial z_{1t}}-\frac{\partial^{2}\eta_{1t}(u_{0})}{\partial z_{1t}},...,\frac{\partial^{2}\eta_{nt}(u_{r})}{\partial z_{nt}}-\frac{\partial^{2}\eta_{nt}(u_{0})}{\partial z_{nt}},$ (A23)
$\frac{\partial^{3}\eta_{1t}(u_{r})}{\partial z_{1t}^{2}}-\frac{\partial^{3}\eta_{1t}(u_{0})}{\partial z_{1t}^{2}},...,\frac{\partial^{3}\eta_{nt}(u_{r})}{\partial z_{nt}^{2}}^{\intercal}-\frac{\partial^{3}\eta_{nt}(u_{0})}{\partial z_{nt}^{2}}\Big{)}^{\intercal}.$
@@
As provided below, in our case, the identifiability of $\mathbf{z}_{t}$ is guaranteed by the linear independence of the whole function vectors $\mathbf{s}_{lt}(u_{r})$ and $\mathring{\mathbf{s}}_{t}(u_{r})$, with $l=1,2,...,n$ and every $u_{r}$. This linear independence is generally a much stronger condition. Theorem A1 can be considered as a special case where the number of domains $u_{r}$ is $1$. In this case, only $\mathbf{s}_{lt}(u_{0})$ in Eq A22 is utilized but $2n$ values of $z_{l,t-r-1}$ are required. Otherwise, in the nonstationary case, the domain information $u_{r}$ can increase the changeability of $s_{lt}(u_{r})$. Besides, $\mathring{\mathbf{s}}_{t}(u_{r})$ in Eq A23 can also help to find more independent vectors to satisfy the sufficiency assumption.
Corollary A1 (Identifiability under Non-Stationary Process).
@@
Suppose $\mathbf{x}_{t}=\mathbf{g}(\mathbf{z}_{t:t-r})$, $\mathbf{z}_{t}=\mathbf{m}(\mathbf{x}_{t:t-\mu})$, and that the conditional distribution $p(z_{kt}\,|\,\mathbf{z}_{t-1:t-r-1},\mathbf{u})$ may change across $a+1$ values of the auxiliary variable $\mathbf{u}$, denoted by $u_{0}$, $u_{1}$, …, $u_{a}$. Suppose the components of $\mathbf{z}_{t}$ are mutually independent conditional on $\mathbf{z}_{t-1:t-r-1}$ with each auxiliary variable. Assume that the components of $\hat{\mathbf{z}}_{t}$ are also mutually independent conditional on $\hat{\mathbf{z}}_{t-1:t-r-1}$. Suppose the domain is path-connected and $\mathbf{m},\mathbf{\hat{m}},\mathbf{g},\mathbf{\hat{g}}$ are second-order differentiable and their combination subjects to non-degenerate condition. If there exists $2n$ different values of function vectors $\mathbf{s}_{lt}(u_{r})$ or $\mathring{\mathbf{s}}_{t}(u_{r})$ and $\mathring{\mathbf{s}}_{t}(u_{r})$, with $l=1,2,...,n$ and every $u_{r}$, are linearly independent, then $\hat{\mathbf{z}}_{t}$ is a permuted invertible component-wise transformation of $\mathbf{z}_{t}$.